OpenStudy (anonymous):

-√3(√-3) help me pleasee

OpenStudy (anonymous):

\(\sqrt{-3}=\sqrt{3}i\)

OpenStudy (anonymous):

i know this is with i agai

OpenStudy (anonymous):

so \[-\sqrt{3}\times \sqrt{3}i=-3i\]

OpenStudy (anonymous):

ohh so anything with a negative is and imaginary number right

OpenStudy (anonymous):

if you have a negative number inside the radical, then it comes outside as \(i\)

OpenStudy (anonymous):

yeah because a negative isnt possible in square root

OpenStudy (anonymous):

for example \(\sqrt{-25}=\sqrt{25}i=5i\) and \[\sqrt{-3}=\sqrt{3}i\]

OpenStudy (anonymous):

i have a few more question like this can i show you and can you tell me if im doing it right?

OpenStudy (anonymous):

√-25√-9 its going to be 5i 3i right?

OpenStudy (anonymous):

yes, that would be the first step

OpenStudy (anonymous):

the second step would be to actually multiply and get \[5i\times 3i=15i^2=15\times (-1)=-15\]

OpenStudy (anonymous):

the i goes away?

OpenStudy (anonymous):

yes, because \(i=\sqrt{-1}\) and so \(i^2=-1\)

OpenStudy (anonymous):

oh that makes sense and im sorry i just have two more if u could please check them as well (6i)2 which would be 36i right but if i multiplied 36 by -1 then i get -36 then theres (2+3)(2-3i) im not sure about this one like is it going to be like 5*2-3(-1) ?

OpenStudy (anonymous):

thanks for helping me out im stuck with all this and guys are the best :D